Last year Ana spent the following:
30% of her salary on rent
2\5 of her salary on entertainment
1\4 of her salary on living expenses.
Ana spent 7200 US Dollars in total.
She saved the rest of her salary.
Work out how much money she saved.

Found 2 solutions by Alan3354, skeet:
Answer by Alan3354(69443)

Last year Ana spent the following:
30% of her salary on rent
2\5 of her salary on entertainment
1\4 of her salary on living expenses.
Ana spent 7200 US Dollars in total.
She saved the rest of her salary.
Work out how much money she saved.
========================
Change it all to fractions.
3/10 of her salary on rent
2/5 of her salary on entertainment
1/4 of her salary on living expenses.
Add them
3/10 + 2/5 + 1/4 = 6/20 + 8/20 + 5/20 = 19/20*S
(19/20)*S = 7200
S = 7200*20/19
------
She saved the remaining 1/20 of S = 7200/19
saved =~ $378.95

Answer by skeet(1)